Foundation November 2020 Paper 1 Q10
10. An art shop gives away a free copy of a photograph with purchases over £10.
The table shows the probability that each photograph, chosen at random, is given away.
| Photograph | Flower | Mountain | Water | City |
|---|---|---|---|---|
| Probability | 0·32 | 0·28 | 0·25 | 0·15 |
(a) Copies of these 4 photographs are the only photographs given away by the art shop in this offer.
Explain how you know this. [1]
Explain how you know this. [1]
(b) Work out the probability that the photograph given away by the art shop is of the Water or City. [1]
| Scheme | Marks | Comments |
|---|---|---|
| 0.32 + 0.28 + 0.25 + 0.15 = 1 (and valid conclusion e.g. ‘so this is certain.’) | E1 | Allow e.g. ‘They all add up to 1’; must be 1 or 1.0 or 1.00; if 100% seen must also see 32% etc to earn the mark Do not allow ‘We have been told there are 4 pictures.’ |
| Scheme | Marks | Comments |
|---|---|---|
| 0.4(0) oe | B1 | Must be a single value; ignore subsequent conversion to different form |
| (2) |